A sample is observed to exert a pressure of 600 torr when confirmed to a 10.0 L vessels at 27 . C what will bbe its pressure in kPa when the gas is transfereed to an 8.0 L vessels at 300K?


1
Expert's answer
2021-05-25T07:49:37-0400

"P_1 = 600 \;torr \\

V_1 = 10.0 \;L \\

T_1 = 27 \;C = 27 + 273 = 300 \;K \\

V_2 = 8.0 \;L \\

T_2 = 300 \;K \\

T_1 = T_2 = constant \\

P_1V_1 = P_2V_2 \\

P_2 = \frac{P_1V_1}{V_2} \\

= \frac{600 \times 10.0}{8.0}= 750 \;torr = 100 \;kPa"

Answer: 100 kPa